# Topological regularity of Busemann spaces of nonpositive curvature

### v1

@@ 5.2. Fibration property
The following two theorems are the main results of this subsection; compare with [64, Theorem 1.11] and [65, Theorem 5.1], respectively. Let $X$ be a GNPC space, and let $\delta_{k}$ be a positive constant depending only on $k$, as in Notation 5.8.
@@
The following is a criterion for the $\varepsilon$-openness (cf. [80, 2.2.1], [64, Lemma 8.1]; see [63, Theorem 1.2] for a more general statement). We say that a metric space $Y$ is locally geodesic if each point has a neighborhood $U$ that is geodesic in $Y$, i.e., any two points in $U$ is connected by a shortest path in $Y$.
@@
We now prove the openness of a strainer map, which is the first half of Theorem 5.10. Compare the statement below with [64, Lemma 8.2].
@@
The proof is similar to that of [64, Lemma 8.2], except that we must deal with the asymmetry of a strainer. Given $v\in\mathbb{R}^{k}$ near $f(x)$, we move $x$ toward the strainer $p_{1},\dots,p_{k}$ or the opposite strainer $\bar{p}_{1},\dots,\bar{p}_{k}$ so that the image of $f$ approaches to the desired value $v$. Here, we modify the norm of $\mathbb{R}^{k}$ so that the image of $f$ approaches to $v$ at a constant speed with respect to this new norm. Since our definition of a strainer is inductive, we use induction on $k$.
@@
To this end, we introduce a new norm on $\mathbb{R}^{k}$. For $v=(v_{[k-1]},v)\in\mathbb{R}^{k-1}\times\mathbb{R}$, we define
@@
We next prove the local uniform contractibility of the fibers of a strainer map, the second half of Theorem 5.10. Since a GNPC space is locally contractible, it suffices to construct a local uniform retraction to the fiber of a strainer map. The construction is a refinement of the previous proof of the openness. Here, we do require the almost nonbranching property, in addition to the almost orthogonality (cf. Remark 5.18). We also need the notion of straining radius, Definition 5.3, to guarantee the uniformity of the retraction. Compare the following with [64, Theorem 9.1], and see also Remarks 5.22 and 5.23.
@@
The proof is similar to that of [64, Theorem 9.1]. The desired retraction $\Phi$ is obtained by composing $\Phi_{p_{1}},\dots,\Phi_{p_{k}}$ and $\Phi_{\bar{p}_{1}},\dots,\Phi_{\bar{p}_{k}}$ infinitely many times, where $\bar{p}_{1},\dots,\bar{p}_{k}$ is an opposite strainer. As before, we have to take into account the asymmetry of the almost orthogonality. We use induction on $k$.
@@
As can be seen from the above proof, the retraction $\Phi$ is actually given by a homotopy as in the original statement [64, Theorem 9.1]: there exists a continuous map
@@
Moreover, the following refinement of the inequality (5.10) holds (compare with [64, Theorem 9.1(2)]):
@@
On the other hand, unlike the GCBA case, it seems impossible to expect that the image of the retraction $\Phi$ remains in the original ball $B(x,r)$ (compare with [64, Theorem 9.1(3)]). We deal with this problem by complementing a strainer in the proof of Theorem 5.11 below. See also Corollary 5.27.
@@ 5.3. Consequences
In the GCBA or Alexandrov case, $f$ is a $\varkappa(\delta)$-almost isometry to a Euclidean domain, i.e., a bi-Lipschitz homeomorphism with Lipschitz constants $1\pm\varkappa(\delta)$ ([64, Corollary 11.2], [21, Theorem 9.4]). Since our GNPC spaces are modeled on (strictly convex) normed spaces, we cannot expect such an improvement of the Lipschitz constants. However, see also Problems 8.3, 8.4, and 8.5.
@@
The above theorem is still a little weaker than the dichotomy stated in [64, Section 9.2, Corollary 11.2]. The technical reason behind this difference is that the closeness of $x$ to $q$ in Proposition 4.14 depends on $q$. In the GCBA (or Alexandrov) case, if $p$ is a $1$-strainer at $q$, then it depends only on the straining radius, independent of $q$ (cf. Remark 5.4). This also seems related to the above-mentioned problems in Section 8.
@@ 6. Extended strainers
In this section, we modify the definition of a strainer to introduce an extended strainer (cf. [66, Section 5]). Roughly speaking, we add a base point $p_{0}$ to a strainer $p_{1},\dots,p_{k}$ such that $p_{1},\dots,p_{k}$ are contained in a metric sphere centered at $p_{0}$, which may be arbitrarily large. This modification is necessary for dealing with a global problem like Theorem 1.3.
@@ 6.1. Definition and existence
We show the existence of an extended strainer. For convenience, we say that $p\in B$ is an extended $(0,\delta)$-strainer for any $x\in B\setminus\{p\}$ and $\delta>0$ (there is no condition, and thus it exists). Compare the following with [66, Lemma 5.3].
@@ 6.2. Fibration property
The underlying idea in the proofs is straightforward: we use the retraction onto $\bar{B}(p_{0},R)$ along the shortest paths to $p_{0}$ so that the desired properties of $f$ still hold when restricted to $\bar{B}(p_{0},R)$. However, unlike the GCBA case [66], this retraction is not “almost orthogonal” to the strainer $p_{1},\dots,p_{k}$, due to the asymmetry issue. More precisely, if we move $x$ toward $p_{0}$, the distances from $p_{1},\dots,p_{k}$ may change significantly. Instead, by Definition 6.1(2), $p_{1},\dots,p_{k}$ are almost orthogonal to this retraction, in the sense that moving $x$ toward $p_{1},\dots,p_{k}$ or its opposite strainer $\bar{p}_{1},\dots,\bar{p}_{k}$ leaves the distance from $p_{0}$ almost unchanged. This is enough to prove the desired properties of the half strainer map $f_{\bar{B}(p_{0},R)}$.
@@
Unlike Theorems 5.10 and 5.11, we cannot find exact counterparts to the above theorems in the original paper of Lytchak–Nagano–Stadler [66] (partial counterparts are [66, Lemma 5.1, Proposition 7.1]). These are another small contribution of this paper. Later we derive from the above theorems the same propositions as in the original paper: see Propositions 6.18, 6.21, and 7.6.
We first prove Theorem 6.7 along the same lines as Theorem 5.10. The following is a counterpart to Proposition 5.17. Compare with [66, Lemma 5.1].
@@
The following is a counterpart to Proposition 5.19. Compare with [66, Proposition 7.1]
@@ 6.3. Consequences
In this subsection, using the properties of extended strainer maps, we prove two propositions which will be used in the proof of Theorem 1.3. These correspond to [66, Proposition 5.5, Corollary 7.2].
The first is the contractibility of a “hemisphere” in the fiber of a half strainer map, which corresponds to [66, Proposition 5.5]. The proof is somewhat different from the original which used a limiting argument with the property of the tangent cone of a GCBA space. Here we give in some sense a more direct proof.
@@
Note that $\Pi_{+}\cap\partial B(x,r)$, as well as $\Pi_{+}(v)\cap\partial B(x,r)$, is locally contractible, as already seen in the proof of Corollary 6.17 (compare with the original statement of [66, Proposition 5.5]). Proposition 6.18 is still a little weaker than [66, Proposition 5.5] because the original statement perturbs not only the value of $f$ but also the distance $R$ from $p_{0}$. However, this simpler statement is enough for the proof of Theorem 7.5, as it takes place on a fixed metric sphere $\partial B(p_{0},R)$.
@@
The second is the homotopy negligibleness of the “boundary” of the fiber of a half strainer map, which corresponds to [66, Corollary 7.2]. In fact, we do not need this because its consequence, Proposition 7.6, is now deduced more directly from Corollaries 6.17 and 6.20 (cf. Remark 7.7). Nevertheless, for comparison we include the statement and proof.
@@
The proof is along basically the same lines as [66, Corollary 7.2], but requires an additional argument to handle the problem mentioned in Remark 5.23. Let $U$ be an open ball around $x$ of the straining radius of the extended strainer defining $f|_{\bar{B}(p_{0},R)}$. By Corollaries 5.24 and 6.15, $\Pi_{+}\cap U$ is locally contractible, and hence an ANR.
@@ 7.1. Local results
The proofs of Theorems 1.1 and 1.2 are based on the following technical result, which is a BNPC counterpart to [65, Theorem 6.3] of Lytchak–Nagano. Recall that a $(k,\delta)$-strainer map $f$ at $x\in X$ is complementable at $x$ if there exists $p\in X$ such that $(f,d(p,\cdot))$ is a $(k+1,\delta)$-strainer map at $x$ (Definition 5.28). For convenience, we treat the constant map to $\mathbb{R}^{0}$ as a $(0,\delta)$-strainer map for any $\delta>0$.
@@
The proof is exactly the same as that of [65, Theorem 6.3]. Here we only explain why the original proof works in our setting.
In the original CAT(00) case, the proof of this theorem depends only on the properties (0)–(8) for strainer maps listed in [65, Section 5]. In fact, the definition of a strainer was not even stated in [65], which means that the arguments there do not rely on the specific definition of a strainer. Therefore it is sufficient to check that all these properties hold for our GNPC strainer map. In what follows we refer to the original paper [65].
Properties (0)–(4) are trivial. Properties (5) and (6) were proved in Propositions 5.5 and 5.7, respectively. Property (7) follows from the extension property (6), Proposition 5.7 (the “stronger form” [64, Theorem 10.5] mentioned there is not essential). Regarding Property (8), we proved the openness in Theorem 5.10 (Proposition 5.17). A slight difference is that, as noted in Remark 5.23, it is unclear in the GNPC case whether a metric ball of straining radius in the fiber of a strainer map is contractible in itself (compare with Corollary 5.27 where the radius depends on the center). Still, the required fibration property stated in [65, Theorem 5.1] holds, as we proved in Theorem 5.11 and Corollary 5.25.
Therefore the original argument from [65, Theorem 6.3] carries over verbatim. The set $E$ is finite (Property (7)). The fact that $\Pi\setminus E$ is a topological $k$-manifold is proved by induction on $k$. The $k=0$ case follows from Theorem 5.30, but actually we do not need this for induction. The $k=1,2$ cases follow from Theorems 2.10 and 2.14 and the fibration property of strainer maps. For $k\geq 3$, we need the inductive assumption (hence the actual base case is $k=2$). The $k=3,4$ cases follow from the the fibration property and the fiber bundle recognition ([65, Section 4.3]). For $k\geq 5$, the proof is much more complicated, which utilizes the disjoint disk property ([65, Section 6.1]) via the fibration property. We omit these technical details. ∎
The $k=n=3$ case gives an alternative proof of the first half of Theorem 3.8, originally by Thurston (cf. the first half of [65, Theorem 6.4]).
@@
Step 2. We next show the “if” part. Suppose for any $p\in X$, any sufficiently small punctured ball around $p$ is homotopy equivalent to $S^{n-1}$. Since a small metric ball is contractible, this implies that $X$ is an ANR homology $n$-manifold (by the excision theorem and the Mayer–Vietoris exact sequence). By Theorem 1.2 proved above, the singular points of $X$ are discrete. Furthermore, by Corollary 7.2, we may assume $n\geq 4$. Now the claim follows from the manifold recognition [65, Theorem 6.2], as $S^{n-1}$ is simply connected. ∎
The following is a BNPC counterpart to the second half of [65, Theorem 6.4].
@@
As mentioned in Section 1.2, the above result does not hold for dimension $\geq 5$, due to the double suspension theorem ([25], [35], cf. [14]). A counterexample is given by the Euclidean cone over the spherical suspension of a CAT($1$) homology sphere. In general we have the following local structure around a singular point. The proof is exactly the same as [65, Theorem 6.5]. The $k=5$ case is attributed to Steve Ferry.
@@ 7.2. Global results
In this subsection, we prove global results on BNPC manifolds, including Theorem 1.3. The main result is the following BNPC counterpart to [66, Theorem 1.4] of Lytchak–Nagano–Stadler.
@@
The proof of Theorem 7.5 is along the same lines with the original proof of [66, Theorem 1.4], which is based only on [66, Propositions 5.4, 5.5]. Recall that we have already proved a BNPC counterpart to [66, Proposition 5.5]: see Proposition 6.18 (and Remark 6.19). Hence it remains to prove a BNPC counterpart to [66, Proposition 5.4].
The following is a BNPC counterpart to [66, Proposition 5.4]. The setting is the same as Proposition 6.21, except for the assumption that $X$ is a homology manifold. Again, note that our terminology and notation are slightly different from the original ones: see Remarks 6.6 and 6.12.
@@
Using the properties of half strainer maps, we provide a little more direct proof than the original one of [66, Proposition 5.4] (see Remark 7.7 below).
@@
Since $\hat{\Pi}\cap U$ is homotopy negligible in $\Pi_{+}\cap U$ by Proposition 6.21, the vanishing of the local homology also follows from a result of Toruńczyk [92, Corollary 2.6] (cf. [93]), as in the original proof of [66, Proposition 5.4].
Now we can prove Theorem 7.5 in exactly the same way as the original proof in [66, Section 6].
@@
The proof is exactly the same as the original proof of [66, Theorem 1.4]. The original proof relies only on the existence of extended strainers ([66, Lemma 5.3]) and the properties of the “half spaces” and “hemispheres” associated with extended strainers ([66, Propositions 5.4, 5.5]). The existence of extended strainers was established in Proposition 6.4, and the properties of the half spaces and hemispheres were proved in Propositions 7.6 and 6.18 (cf. Remark 6.19). Therefore the original proof, using the manifold recognition in dimension $\leq 2$ (Theorem 2.10) and the Jordan–Schoenflies theorem, still works in our setting. We leave it to the reader to check the details. ∎
The following corollaries are proved in exactly the same way as [66, Corollaries 1.5, 1.6], only using the geodesic contraction with topological arguments.

### v2

@@ 5.2. Fibration property
The following two theorems are the main results of this subsection; compare with [63, Theorem 1.11] and [64, Theorem 5.1], respectively. Let $X$ be a GNPC space, and let $\delta_{k}$ be a positive constant depending only on $k$, as in Notation 5.8.
@@
The following is a criterion for the $\varepsilon$-openness (cf. [79, 2.2.1], [63, Lemma 8.1]; see [62, Theorem 1.2] for a more general statement). We say that a metric space $Y$ is locally geodesic if each point has a neighborhood $U$ that is geodesic in $Y$, i.e., any two points in $U$ is connected by a shortest path in $Y$.
@@
We now prove the openness of a strainer map, which is the first half of Theorem 5.10. Compare the statement below with [63, Lemma 8.2].
@@
The proof is similar to that of [63, Lemma 8.2], except that we must deal with the asymmetry of a strainer. Given $v\in\mathbb{R}^{k}$ near $f(x)$, we move $x$ toward the strainer $p_{1},\dots,p_{k}$ or the opposite strainer $\bar{p}_{1},\dots,\bar{p}_{k}$ so that the image of $f$ approaches to the desired value $v$. Here, we modify the norm of $\mathbb{R}^{k}$ so that the image of $f$ approaches to $v$ at a constant speed with respect to this new norm. Since our definition of a strainer is inductive, we use induction on $k$.
@@
To this end, we introduce a new norm on $\mathbb{R}^{k}$. For $v=(v_{[k-1]},v_{k})\in\mathbb{R}^{k-1}\times\mathbb{R}$, we define
@@
We next prove the local uniform contractibility of the fibers of a strainer map, the second half of Theorem 5.10. Since a GNPC space is locally contractible, it suffices to construct a local uniform retraction to the fiber of a strainer map. The construction is a refinement of the previous proof of the openness. Here, we do require the almost nonbranching property, in addition to the almost orthogonality (cf. Remark 5.18). We also need the notion of straining radius, Definition 5.3, to guarantee the uniformity of the retraction. Compare the following with [63, Theorem 9.1], and see also Remarks 5.22 and 5.23.
@@
The proof is similar to that of [63, Theorem 9.1]. The desired retraction $\Phi$ is obtained by composing $\Phi_{p_{1}},\dots,\Phi_{p_{k}}$ and $\Phi_{\bar{p}_{1}},\dots,\Phi_{\bar{p}_{k}}$ infinitely many times, where $\bar{p}_{1},\dots,\bar{p}_{k}$ is an opposite strainer. As before, we have to take into account the asymmetry of the almost orthogonality. We use induction on $k$.
@@
As can be seen from the above proof, the retraction $\Phi$ is actually given by a homotopy as in the original statement [63, Theorem 9.1]: there exists a continuous map
@@
Moreover, the following refinement of the inequality (5.10) holds (compare with [63, Theorem 9.1(2)]):
@@
On the other hand, unlike the GCBA case, it seems impossible to expect that the image of the retraction $\Phi$ remains in the original ball $B(x,r)$ (compare with [63, Theorem 9.1(3)]). We deal with this problem by complementing a strainer in the proof of Theorem 5.11 below. See also Corollary 5.27.
@@ 5.3. Consequences
In the GCBA or Alexandrov case, $f$ is a $\varkappa(\delta)$-almost isometry to a Euclidean domain, i.e., a bi-Lipschitz homeomorphism with Lipschitz constants $1\pm\varkappa(\delta)$ ([63, Corollary 11.2], [21, Theorem 9.4]). Since our GNPC spaces are modeled on (strictly convex) normed spaces, we cannot expect such an improvement of the Lipschitz constants. However, see also Problems 8.3, 8.4, and 8.5.
@@
The above theorem is still a little weaker than the dichotomy stated in [63, Section 9.2, Corollary 11.2]. The technical reason behind this difference is that the closeness of $x$ to $q$ in Proposition 4.14 depends on $q$. In the GCBA (or Alexandrov) case, if $p$ is a $1$-strainer at $q$, then it depends only on the straining radius, independent of $q$ (cf. Remark 5.4). This also seems related to the above-mentioned problems in Section 8.
@@ 6. Extended strainers
In this section, we modify the definition of a strainer to introduce an extended strainer (cf. [65, Section 5]). Roughly speaking, we add a base point $p_{0}$ to a strainer $p_{1},\dots,p_{k}$ such that $p_{1},\dots,p_{k}$ are contained in a metric sphere centered at $p_{0}$, which may be arbitrarily large. This modification is necessary for dealing with a global problem like Theorem 1.3.
@@ 6.1. Definition and existence
We show the existence of an extended strainer. For convenience, we say that $p\in B$ is an extended $(0,\delta)$-strainer for any $x\in B\setminus\{p\}$ and $\delta>0$ (there is no condition, and thus it exists). Compare the following with [65, Lemma 5.3].
@@ 6.2. Fibration property
The underlying idea in the proofs is straightforward: we use the retraction onto $\bar{B}(p_{0},R)$ along the shortest paths to $p_{0}$ so that the desired properties of $f$ still hold when restricted to $\bar{B}(p_{0},R)$. However, unlike the GCBA case [65], this retraction is not “almost orthogonal” to the strainer $p_{1},\dots,p_{k}$, due to the asymmetry issue. More precisely, if we move $x$ toward $p_{0}$, the distances from $p_{1},\dots,p_{k}$ may change significantly. Instead, by Definition 6.1(2), $p_{1},\dots,p_{k}$ are almost orthogonal to this retraction, in the sense that moving $x$ toward $p_{1},\dots,p_{k}$ or its opposite strainer $\bar{p}_{1},\dots,\bar{p}_{k}$ leaves the distance from $p_{0}$ almost unchanged. This is enough to prove the desired properties of the half strainer map $f|_{\bar{B}(p_{0},R)}$.
@@
Unlike Theorems 5.10 and 5.11, we cannot find exact counterparts to the above theorems in the original paper of Lytchak–Nagano–Stadler [65] (partial counterparts are [65, Lemma 5.1, Proposition 7.1]). These are another small contribution of this paper. Later we derive from the above theorems the same propositions as in the original paper: see Propositions 6.18, 6.21, and 7.6.
We first prove Theorem 6.7 along the same lines as Theorem 5.10. The following is a counterpart to Proposition 5.17. Compare with [65, Lemma 5.1].
@@
The following is a counterpart to Proposition 5.19. Compare with [65, Proposition 7.1]
@@ 6.3. Consequences
In this subsection, using the properties of extended strainer maps, we prove two propositions which will be used in the proof of Theorem 1.3. These correspond to [65, Proposition 5.5, Corollary 7.2].
The first is the contractibility of a “hemisphere” in the fiber of a half strainer map, which corresponds to [65, Proposition 5.5]. The proof is somewhat different from the original which used a limiting argument with the property of the tangent cone of a GCBA space. Here we give in some sense a more direct proof.
@@
Note that $\Pi_{+}\cap\partial B(x,r)$, as well as $\Pi_{+}(v)\cap\partial B(x,r)$, is locally contractible, as already seen in the proof of Corollary 6.17 (compare with the original statement of [65, Proposition 5.5]). Proposition 6.18 is still a little weaker than [65, Proposition 5.5] because the original statement perturbs not only the value of $f$ but also the distance $R$ from $p_{0}$. However, this simpler statement is enough for the proof of Theorem 7.5, as it takes place on a fixed metric sphere $\partial B(p_{0},R)$.
@@
The second is the homotopy negligibleness of the “boundary” of the fiber of a half strainer map, which corresponds to [65, Corollary 7.2]. In fact, we do not need this because its consequence, Proposition 7.6, is now deduced more directly from Corollaries 6.17 and 6.20 (cf. Remark 7.7). Nevertheless, for comparison we include the statement and proof.
@@
The proof is along basically the same lines as [65, Corollary 7.2], but requires an additional argument to handle the problem mentioned in Remark 5.23. Let $U$ be an open ball around $x$ of the straining radius of the extended strainer defining $f|_{\bar{B}(p_{0},R)}$. By Corollaries 5.24 and 6.15, $\Pi_{+}\cap U$ is locally contractible, and hence an ANR.
@@ 7.1. Local results
The proofs of Theorems 1.1 and 1.2 are based on the following technical result, which is a BNPC counterpart to [64, Theorem 6.3] of Lytchak–Nagano. Recall that a $(k,\delta)$-strainer map $f$ at $x\in X$ is complementable at $x$ if there exists $p\in X$ such that $(f,d(p,\cdot))$ is a $(k+1,\delta)$-strainer map at $x$ (Definition 5.28). For convenience, we treat the constant map to $\mathbb{R}^{0}$ as a $(0,\delta)$-strainer map for any $\delta>0$.
@@
The proof is exactly the same as that of [64, Theorem 6.3]. Here we only explain why the original proof works in our setting.
In the original CAT(00) case, the proof of this theorem depends only on the properties (0)–(8) for strainer maps listed in [64, Section 5]. In fact, the definition of a strainer was not even stated in [64], which means that the arguments there do not rely on the specific definition of a strainer. Therefore it is sufficient to check that all these properties hold for our GNPC strainer map. In what follows we refer to the original paper [64].
Properties (0)–(4) are trivial. Properties (5) and (6) were proved in Propositions 5.5 and 5.7, respectively. Property (7) follows from the extension property (6), Proposition 5.7 (the “stronger form” [63, Theorem 10.5] mentioned there is not essential). Regarding Property (8), we proved the openness in Theorem 5.10 (Proposition 5.17). A slight difference is that, as noted in Remark 5.23, it is unclear in the GNPC case whether a metric ball of straining radius in the fiber of a strainer map is contractible in itself (compare with Corollary 5.27 where the radius depends on the center). Still, the required fibration property stated in [64, Theorem 5.1] holds, as we proved in Theorem 5.11 and Corollary 5.25.
Therefore the original argument from [64, Theorem 6.3] carries over verbatim. The set $E$ is finite (Property (7)). The fact that $\Pi\setminus E$ is a topological $k$-manifold is proved by induction on $k$. The $k=0$ case follows from Theorem 5.30, but actually we do not need this for induction. The $k=1,2$ cases follow from Theorems 2.10 and 2.14 and the fibration property of strainer maps. For $k\geq 3$, we need the inductive assumption (hence the actual base case is $k=2$). The $k=3,4$ cases follow from the the fibration property and the fiber bundle recognition ([64, Section 4.3]). For $k\geq 5$, the proof is much more complicated, which utilizes the disjoint disk property ([64, Section 6.1]) via the fibration property. We omit these technical details. ∎
The $k=n=3$ case gives an alternative proof of the first half of Theorem 3.9, originally by Thurston (cf. the first half of [64, Theorem 6.4]).
@@
Step 2. We next show the “if” part. Suppose for any $p\in X$, any sufficiently small punctured ball around $p$ is homotopy equivalent to $S^{n-1}$. Since a small metric ball is contractible, this implies that $X$ is an ANR homology $n$-manifold (by the excision theorem and the Mayer–Vietoris exact sequence). By Theorem 1.2 proved above, the singular points of $X$ are discrete. Furthermore, by Corollary 7.2, we may assume $n\geq 4$. Now the claim follows from the manifold recognition [64, Theorem 6.2], as $S^{n-1}$ is simply connected. ∎
The following is a BNPC counterpart to the second half of [64, Theorem 6.4].
@@
As mentioned in Section 1.2, the above result does not hold for dimension $\geq 5$, due to the double suspension theorem ([25], [35], cf. [14]). A counterexample is given by the Euclidean cone over the spherical suspension of a CAT($1$) homology sphere. In general we have the following local structure around a singular point. The proof is exactly the same as [64, Theorem 6.5]. The $k=5$ case is attributed to Steve Ferry.
@@ 7.2. Global results
In this subsection, we prove global results on BNPC manifolds, including Theorem 1.3. The main result is the following BNPC counterpart to [65, Theorem 1.4] of Lytchak–Nagano–Stadler.
@@
The proof of Theorem 7.5 is along the same lines with the original proof of [65, Theorem 1.4], which is based only on [65, Propositions 5.4, 5.5]. Recall that we have already proved a BNPC counterpart to [65, Proposition 5.5]: see Proposition 6.18 (and Remark 6.19). Hence it remains to prove a BNPC counterpart to [65, Proposition 5.4].
The following is a BNPC counterpart to [65, Proposition 5.4]. The setting is the same as Proposition 6.21, except for the assumption that $X$ is a homology manifold. Again, note that our terminology and notation are slightly different from the original ones: see Remarks 6.6 and 6.12.
@@
Using the properties of half strainer maps, we provide a little more direct proof than the original one of [65, Proposition 5.4] (see Remark 7.7 below).
@@
Since $\hat{\Pi}\cap U$ is homotopy negligible in $\Pi_{+}\cap U$ by Proposition 6.21, the vanishing of the local homology also follows from a result of Toruńczyk [92, Corollary 2.6] (cf. [93]), as in the original proof of [65, Proposition 5.4].
Now we can prove Theorem 7.5 in exactly the same way as the original proof in [65, Section 6].
@@
The proof is exactly the same as the original proof of [65, Theorem 1.4]. The original proof relies only on the existence of extended strainers ([65, Lemma 5.3]) and the properties of the “half spaces” and “hemispheres” associated with extended strainers ([65, Propositions 5.4, 5.5]). The existence of extended strainers was established in Proposition 6.4, and the properties of the half spaces and hemispheres were proved in Propositions 7.6 and 6.18 (cf. Remark 6.19). Therefore the original proof, using the manifold recognition in dimension $\leq 2$ (Theorem 2.10) and the Jordan–Schoenflies theorem, still works in our setting. We leave it to the reader to check the details. ∎
The following corollaries are proved in exactly the same way as [65, Corollaries 1.5, 1.6], only using the geodesic contraction with topological arguments.